OpenStudy (anonymous):

Evaluate a-2 for a = 8. I have no clue what it is..

OpenStudy (debbieg):

A variable is just a "placeholder" for a number here. So you just "plug in" the value for a into the expression, and then simplify it. I'll do a different one as an example: Evaluate z+5 for z=3 So I put in a "3" in place of the "z": 3+5=8 Got it? You try it for your problem now. :)

OpenStudy (anonymous):

My options were 1/-64 1/-16 1/16 and –64 and I got none of them, I still don't get it.

OpenStudy (amistre64):

then youve most likely omited the exponent ^ part

OpenStudy (anonymous):

Oh, is it -64?

OpenStudy (amistre64):

no, but it does have a 64 in it

OpenStudy (amistre64):

\[8^{-2}=\frac{1}{8^2}\]

OpenStudy (anonymous):

it's 1/-64? right?

OpenStudy (amistre64):

there would be no negative in it ... which makes me wonder about these options

OpenStudy (amistre64):

i would assume thats the "most correct" option tho .. assuming that youve left out some of it in the post

OpenStudy (anonymous):

oh sorry it's 1/64.. That was another option.

OpenStudy (amistre64):

:) well, id go with the 1/64 then

OpenStudy (debbieg):

lol... yeah, knowing that the -2 was a power would have been useful. :) If you don't like the equation editor, just use a ^ and ( ) as necessary: a^(-2) is unambiguous, and means something much different from a-2 :)